# Nuanced Safety in Generative AI: How Demographics Shape Responsiveness to Severity

### v1

@@ Abstract
Ensuring safety of Generative AI requires a nuanced understanding of pluralistic viewpoints. In this paper, we introduce a novel data-driven approach for calibrating granular ratings in pluralistic datasets. Specifically, we address the challenge of interpreting responses of a diverse population to safety expressed via ordinal scales (e.g., Likert scale). We distill non-parametric responsiveness metrics that quantify the consistency of raters in scoring the varying levels of the severity of safety violations. Using safety evaluation of AI-generated content as a case study, we investigate how raters from different demographic groups (age, gender, ethnicity) use an ordinal scale to express their perception of the severity of violations in a pluralistic safety dataset. We apply our metrics across violation types, demonstrating their utility in extracting nuanced insights that are crucial for developing reliable AI systems in a multi-cultural contexts. We show that our approach offers improved capabilities for prioritizing safety concerns by capturing nuanced viewpoints across different demographic groups, hence improving the reliability of pluralistic data collection and in turn contributing to more robust AI evaluations.
Diverse crowd raters, AI Safety, Human preferences, Severity of violations, Generative AI
@@ 1 Introduction
Ensuring safety of Generative AI is paramount for their responsible deployment and societal trust. Recent research demonstrates that safety perceptions are not uniform but vary significantly across individuals and groups (Aroyo et al., 2024; Kirk et al., 2024; Rastogi et al., 2024).
Safety evaluation tasks often use binary ratings, such as safe and unsafe, which lack the granularity needed for effective alignment with human perception (Wu et al., 2023; Collins et al., 2024). To capture more fine-grained perceptions of the severity of perceived harm, safety evaluation tasks also employ ordinal scales, such as Likert scale (Curry et al., 2021), which allow for more nuanced feedback critical in creation of pluralistic datasets. However, such scales are susceptible to increased noise from variations in individual interpretation and usage, including response biases (Paulhus, 1991) such as extreme responses (Greenleaf, 1992), central tendency (for odd scales), and forced choice and polarization (for even scales). Such scoring biases further propagate in Large Language Models (LLMs). On the one hand, when non-calibrated ratings are used for training of LLMs can cause exaggerated safety behaviors. On the other hand, when LLMs are used as raters in evaluation tasks (Bavaresco et al., 2024; Wang et al., 2024) the non-calibrated ratings can degrade the utility of AI models, and in both cases can lead to downstream harms due to aggregation-based approaches of non-calibrated ratings.
This paper introduces a novel data-driven non-parametric approach for calibrating granular ratings in pluralistic datasets, offering a more robust and nuanced evaluation than traditional approaches. Using safety evaluation of AI-generated content as a case study, we address the critical challenge of understanding how raters from diverse demographic groups interpret and utilize ordinal scales (e.g., Likert scales) when assessing the severity of safety violations. The main contributions of this paper are:
Metrics Development:
We distill robust responsiveness metrics from observed data to interpret the scoring patterns of different rater groups for the varying levels of severity. These metrics allow us to:
@@
    Measure responsiveness to severity: How consistently do different rater groups use a given ordinal scale to express the varying levels of severity of violations?
@@
    Compare responsiveness: Are different rater groups equally responsive to the varying levels of severity?
Application and Validation:
We apply these metrics to a pluralistic safety evaluation dataset, demonstrating their utility in extracting nuanced insights that are crucial for developing AI systems in a multi-cultural context by:
@@
  • Capturing nuanced viewpoints: identifying demographic groups most responsive to severity across violation types, resulting in a granular understanding of pluralistic viewpoints;
  • Prioritizing high-impact items: taking items deemed highly unsafe by groups with with high responsiveness to severity;
  • Improving pluralistic data collection: establishing a reliable and repeatable process for sampling raters and rater groups with high responsiveness to severity.
@@ 2 Related work
Collecting human perspectives on behavior of generative AI models is exceedingly commonplace with growth in its usage in real world tasks. Across literature in AI evaluation, different configurations of human feedback have been studied, with an outsized focus on binary (0/1) human feedback. Recent research (Collins et al., 2024; Arhin et al., 2021; Denton et al., 2021) discusses the limitations of binary feedback in capturing the nuance involved in generative AI evaluation, especially in safety. Wu et al. (2023); Collins et al. (2024) propose fine-grained human feedback encompassing evaluation across multiple attributes and with higher density, yielding improvement in downstream AI tasks via RLHF. Further, Rauh et al. (2024); Jiang et al. (2021) emphasise the importance of measuring extent of harm (severity) in evaluation of algorithms. Another dimension in collecting human feedback relates to the identity of the human providing the feedback. The role of rater identity in their annotation has been discussed extensively in AI evaluation literature (Denton et al., 2021; Arhin et al., 2021; Aroyo et al., 2024; Homan et al., 2023). For developing AI that aligns with human values, Sorensen et al. (2024) show the importance of considering pluralistic viewpoints from a diverse set of raters. Our research builds upon this body of work by specifically examining human feedback collected on a fine-grained (0-4) ordinal scale from raters belonging to different groups with different collective identities.
Calibration of human judgements.
@@
Human judgments elicited as scores on a scale are often miscalibrated, implying that the scores given by people are incomparable due to differences in interpretation and usage of each score (see Griffin & Brenner (2008); Poston (2008) and references therein). Miscalibration in human scores is sometimes addressed through simplifying modeling assumptions about how the miscalibration presents in the data. These modeling assumptions include linear models with additive biases corresponding to rater identity (Bürkner & Vuorre, 2019; Paul, 2011; Barr et al., 2013), models with rater identity-based scale-and-shift biases (Paul, 2011; Roos et al., 2011), mixed-effects models, among others (Wang & Shah, 2019). However, research has shown that issues of human judgment calibration are often more complex, causing significant violations to these simplified assumptions (see Griffin & Brenner (2008) and references therein). In this work, while making minimal assumptions on the nature of miscalibration in human judgments, we provide non-parametric metrics to measure the consistency of raters in reflecting varying levels of severity. Traditional non-parametric metrics like Kendall’s $\tau$ and area under the PR or ROC curves do not capture well the responsiveness to severity. Using a real-world dataset, we surface insights from our proposed metrics into the scoring patterns of different rater groups.
@@ 3 Setup
We consider a general setup with two different rater populations that reflect two contrasting safety evaluation paradigms (Rottger et al., 2022): crowd raters who indicate safety perceptions of a diverse population, and expert raters who follow detailed guidelines for prescribing what counts as safe or unsafe.
@@
    Crowd raters provide a set of pluralistic safety perceptions on an ordinal scale, with each of their ratings representing the perception of a certain rater group. Each rater reviews every item in the given dataset and provides an integer score on a (0-$K$) Likert scale, where 0 is not harmful and $K$ is completely harmful.
@@
    Expert raters strictly follow a set of prescribed rules. For each item in the given dataset, they provide a binary score of 0 (safe) or 1 (unsafe).
@@ 3.1 Data model
To formalize our assumptions regarding the scores given by the individual raters, i.e., individual expert raters and individual crowd raters, we define a data model. Let $R_{ij}$ be the latent severity rating of rater $j$ for prompt-image pair $i$. We assume the following ordinal data model:
$R_{ij}=\mathcal{F}(P_{ij},b_{j},c_{i})$
where $P_{ij}$ is the perceived severity of prompt-image pair $i$ by rater $j$, $b_{j}$ is the rater-specific bias in perception, $c_{i}$ is the prompt-image pair specific bias in perception, and $\mathcal{F}$ is some function over these. Then, the relationship between the latent severity ratings and the scores of the two rater populations can be written as:
  • Expert rater: $S_{ij}=1$ if $R_{ij}>t_{j}$, where $S_{ij}$ is the binary score given by expert rater $j$ to prompt-image pair $i$ and $t_{j}$ is the threshold above which they assign a binary score of 1.
  • Crowd rater: $S_{ij}=k$ if $R_{ij}>t_{jk}$, where $S_{ij}$ is the Likert score given by crowd rater $j$ to prompt-image pair $i$ and $t_{jk}$ is the threshold above which they assign the Likert score $k\in\{0,1,\dots,K\}$.
This data model assumes that the perception of severity varies from rater to rater, be they expert or crowd raters. To simplify the data model, we introduce a latent variable $V$ to represent the true, but unobservable, severity of prompt-image pairs. This acknowledges that true severity isn’t directly measurable but is still the underlying factor that monotonically influences every rater’s judgments. In the case of expert raters, $V$ represents the shared understanding of severity as defined by the strict guidelines that they must adhere to, where the guidelines serve as a framework to operationalize the theoretical notion of severity (cf. prescriptive annotation paradigm). In the case of crowd raters, $V$ represents some shared perception of severity based on lived experiences (cf. descriptive annotation paradigm). With the simplification, we can reformulate $R_{ij}$ as:
$R_{ij}=\mathcal{F}^{\prime}(V_{i},b_{j})$
where $V_{i}$ is the true underlying severity of prompt-image pair $i$. This simplification creates a more tractable model without requiring estimation of $P_{ij}$. Moreover, it is akin to the standard assumption of a uni-dimensional latent trait found in methods like Item Response Theory (Samejima, 1968). As such, we work with the simplified data model hereon. However, we note that the simplification limits our ability to exactly compute the component of raters’ individual perceptions in their responsiveness to severity.
@@ 3.2 Responsiveness to severity
Having a Likert scale allows crowd raters to express their safety judgments on a severity spectrum rather than classifying prompt-image pairs as safe or unsafe. However, Likert scales do not guarantee that rater scores will meaningfully reflect the severity of violations. For example, there may be raters who only use the ends of the scale, or those who cluster all their ratings around certain scores. It is essential to disentangle the scale use from the actual response to the severity of violations. While one might ideally want to define responsiveness as a direct relationship between a rater’s scores $S$ and the true severity $V$, this is not practically feasible. True severity $V$ is a theoretical construct that cannot be objectively measured or easily determined. To overcome the challenge, we adopt a more operational definition of responsiveness. We formally define responsiveness to the severity of violations as being composed of the following two properties that can be quantified using observable data:
@@
  • Ability to stochastically order severity. If a rater is responsive to the true underlying severity $V$, then a higher score from them should correspond to a higher probability of the true severity crossing any threshold $T=t$. This is the principle of first-order stochastic dominance, which can be stated as $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ for all $T\in\mathcal{T}$ when $s_{1}>s_{2}$.
@@
  • Ability to discriminate between distinct levels of severity. If a rater is responsive to the true underlying severity $V$, then they should be able to discriminate the prompt-image pairs whose true severities are above a given threshold $T=t$ from those below that threshold. This means that $P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$ for all $T\in\mathcal{T}$.
@@
Intuitively, these two properties together signify that a rater who is responsive to the severity of violations is able to consistently convey the severity of violations at each score of the Likert scale. Our approach to characterizing responsiveness to severity relies on a notion of stochastic ordering that is related to the classic decision-theoretic concepts of stochastic ordering and outcome monotonicity (Birnbaum & Navarrete, 1998). The stochastic ordering property can further be related to monotonicity of a calibration curve or reliability diagram (DeGroot & Fienberg, 1983) in the case that we consider calibration of non-expert scores against some binarized reference for severity. The property concerning the ability of raters to discriminate between distinct levels of severity is related to the notion of discriminability from signal detection theory (McNicol, 2005) used to motivated the design of metrics such as the area under the receiver operating characteristic (ROC) curve and Kendall’s $\tau$ (Kendall, 1938). Furthermore, the property is also related to the notion of discrimination between different levels of the latent trait in Item Response Theory (Samejima, 1968) and Mokken Scale Analysis (Mokken, 1971).
@@ 4 Metric design
Next, our aim is to quantify the responsiveness of the different demographic groups to the severity of violations in order to evaluate and compare them. We achieve such a quantification by individually quantifying the two properties that constitute responsiveness. To do so, we need a signal for $V>T$ since $V$ itself is unobservable. We treat the binary scores of expert raters as binary labels $U$ and take $U=1$ as the signal for $V$ exceeding some $T$. To obtain the binary labels $U$ for the prompt-image pairs, we assign the individual binary scores of every expert to the prompt-image pairs by replicating each prompt-image pair for every expert. Alternatively, we could also obtain the binary labels $U$ for prompt-image pairs from crowd raters themselves, excluding the demographic group being evaluated to maintain the independence of $S$ and $T$: we first binarize the individual Likert scores of the crowd raters using a score $s\in[1,4]$ as boundary, then assign these individual binary scores by replicating each prompt-image pair for every crowd rater.
We note that when the binary labels $U$ are obtained using expert raters, we are quantifying the responsiveness to severity as captured by the expert raters based on the guidelines they use. On the other hand, when the binary labels $U$ are obtained using crowd raters (excluding the group being evaluated), we are quantifying the responsiveness to severity as captured by the collective judgment of the crowd rater population. The latter paradigm is discussed in Section 7.
@@
Here, we reformulate the inequalities presented above that relate to the ability to stochastically order and discriminate between levels of severity to leverage $U=1$ in place of $V>T$. Note that as we assume that the threshold $T$ varies across individual expert raters, it is not straightforward to directly substitute $U=1$ into the threshold-specific inequalities. However, we show that if the two inequalities hold for all individual thresholds $T$, then they also hold for $U=1$.
  • Ability to stochastically order. We had that $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$, when $s_{1}>s_{2}$, for all $T=t$. Hence, when $s_{1}>s_{2}$, $P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$.
  • Ability to discriminate. We had that $P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$ for all $T=t$. Hence, we have $P(S\geq s|U=1)\geq P(S\geq s|U=0)$.
To prove the two properties, let $\mathcal{T}=\{t_{1},\dots,t_{n}\}$ be the set of all thresholds $T$. For the stochastic ordering property, we have $P(U=1|S=s_{1})=\sum_{t\in\mathcal{T}}P(V>T|S=s_{1},T=t)P(T=t)$ and $P(U=1|S=s_{2})=\sum_{t\in\mathcal{T}}P(V>T|S=s_{2},T=t)P(T=t)$ and given that $S$ and $T$ are independent. Since $P(T=t)$ is non-negative and $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ for every $t\in\mathcal{T}$, hence, $P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$.
For the discrimination property, for every $t\in\mathcal{T}$ we have $P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$, which gives $P(S\geq s|U=1,T=t)\geq P(S\geq s|U=0,T=t)$ for any given $t$. By Bayes’ rule, $\frac{P(S\geq s,U=1,T=t)}{P(U=1,T=t)}\geq\frac{P(S\geq s,U=0,T=t)}{P(U=0,T=t)}$ for any given $t$. Let $a(t)=P(S\geq s,U=1,T=t)$, $b(t)=P(U=1,T=t)$, and $c(t)=P(T=t)$. Since $S$ and $T$ are independent, $P(S\geq s,T=t)=P(S\geq s)c(t)$. Then, $P(S\geq s,U=0,T=t)=P(S\geq s)c(t)-a(t)$ and $P(U=0,T=t)=c(t)-b(t)$. So, the inequality becomes $\frac{a(t)}{b(t)}\geq\frac{P(S\geq s)c(t)-a(t)}{c(t)-b(t)}$ for any given $t$. Cross-multiplying and summing the inequalities over all $t\in\mathcal{T}$, we have that $\sum_{t\in\mathcal{T}}a(t)\geq P(S\geq s)\sum_{t\in\mathcal{T}}b(t)$. This yields $P(S\geq s,U=1)\geq P(S\geq s)P(U=1)$, and hence, $P(S\geq s|U=1)\geq P(S\geq s)$. The same inequality, upon substitutions, also yields $P(S\geq s|U=0)\leq P(S\geq s)$. Therefore, $P(S\geq s|U=1)\geq P(S\geq s|U=0)$.

### v4

# Decoding Safety Feedback from Diverse Raters: A Data-driven Lens on Responsiveness to Severity

@@ Abstract
Ensuring the safety of Generative AI requires a nuanced understanding of pluralistic viewpoints. In this paper, we introduce a novel data-driven approach for interpreting granular ratings in pluralistic datasets. Specifically, we address the challenge of analyzing nuanced differences in safety feedback from a diverse population expressed via ordinal scales (e.g., a Likert scale). We distill non-parametric responsiveness metrics that quantify the consistency of raters in scoring varying levels of the severity of safety violations. Leveraging a publicly available pluralistic dataset of safety feedback on AI-generated content as our case study, we investigate how raters from different demographic groups (age, gender, ethnicity) use an ordinal scale to express their perceptions of the severity of violations. We apply our metrics across violation types, demonstrating their utility in extracting nuanced insights that are crucial for aligning AI systems reliably in multi-cultural contexts. We show that our approach can inform rater selection and feedback interpretation by capturing nuanced viewpoints across different demographic groups, hence improving the quality of pluralistic data collection and in turn contributing to more robust AI development.
@@ 1 Introduction
Ensuring the safety of Generative AI is paramount for their responsible deployment and societal trust. Recent research demonstrates that safety perceptions are not uniform but vary significantly across individuals and groups (Aroyo et al., 2024; Kirk et al., 2024; Rastogi et al., 2024).
Safety evaluation tasks often use binary ratings, such as safe and unsafe, which lack the granularity needed for effective alignment with human preferences (Wu et al., 2023; Collins et al., 2024). To capture more fine-grained perceptions of the severity of potential harm, safety alignment tasks increasingly employ ordinal scales, such as Likert scales (Curry et al., 2021), which allow for more nuanced feedback that is critical for the creation of pluralistic datasets for SFT / RLHF. However, these scales are harder to interpret due to variations in individual perceptions and response biases (Paulhus, 1991) such as extreme responses (Greenleaf, 1992), central tendency (for odd scales), and forced choice and polarization (for even scales). Inconsistencies in how raters map perceived severity to ordinal scores can propagate to datasets and optimization objectives for safety alignment, especially in multi-cultural contexts (Kaufmann et al., 2024). This, in turn, can lead to exaggerated safety behaviors in models like LLMs, e.g., false refusal of benign requests or generation of content that is harmful to certain cultures. Moreover, when LLMs are used as judges or reward models (Bavaresco et al., 2024; Wang et al., 2024), such inconsistencies and response biases can degrade their utility. Both cases cause downstream harms due to poor understanding of nuanced safety feedback.
This paper introduces a novel data-driven non-parametric approach for interpreting granular ratings in pluralistic datasets, offering more robust and nuanced insights than traditional approaches. Using safety feedback on AI-generated content as a case study, we address the critical challenge of understanding how raters from diverse demographic groups utilize ordinal scales (e.g., Likert scales) when expressing the severity of safety violations. The main contributions of this paper are:
Metrics development: We distill robust responsiveness metrics from observable data to analyze the feedback of individual raters or rater groups for varying levels of severity. These metrics allow us to:
@@
    Measure responsiveness to severity: How do different raters or rater groups use a given ordinal scale to score the varying levels of severity of violations?
@@
    Compare responsiveness: Do different raters or rater groups respond similarly to varying levels of severity of violations?
Metrics application: We apply these metrics to a publicly available pluralistic dataset of safety feedback on AI-generated content (Rastogi et al., 2024), demonstrating their utility in extracting nuanced insights that are crucial for aligning AI systems in multi-cultural contexts, specifically:
@@
  • Capturing nuanced viewpoints: quantifying the responsiveness of demographic groups to severity across violation types, resulting in a granular understanding of pluralistic viewpoints;
  • Prioritizing high-impact items: taking items deemed highly unsafe by raters with high responsiveness to severity from different groups;
  • Improving pluralistic data collection: establishing a reliable and repeatable process for selecting raters with high responsiveness to severity from different groups.
@@ 2 Related work
Nuanced human feedback for AI safety. Collecting human perspectives on behavior of generative AI models is exceedingly commonplace with growth in its usage in real world tasks. Across the literature on AI evaluation, different configurations of human feedback have been studied, with an outsized focus on binary (0/1) human feedback. Recent research (Collins et al., 2024; Arhin et al., 2021; Denton et al., 2021) discusses the limitations of binary feedback in capturing the nuance involved in generative AI evaluation, especially in safety. Wu et al. (2023); Collins et al. (2024) propose fine-grained human feedback encompassing evaluation across multiple attributes and with higher density, yielding improvement in downstream AI tasks via RLHF. Further, Rauh et al. (2024); Jiang et al. (2021) emphasize the importance of measuring extent of harm (severity) in evaluation of algorithms. Another dimension in collecting human feedback relates to the identity of the human providing the feedback. The role of rater identity in their annotation has been discussed extensively in AI evaluation literature (Denton et al., 2021; Arhin et al., 2021; Aroyo et al., 2024; Homan et al., 2023; Pei & Jurgens, 2023; Davani et al., 2024). For developing AI that aligns with human values, Sorensen et al. (2024) show the importance of considering pluralistic viewpoints from a diverse set of raters. Our research builds upon this body of work by specifically examining human feedback collected on a fine-grained ordinal scale from raters belonging to different groups with different collective identities.
@@
Interpretation and calibration of human judgments. Human judgments elicited as scores on a scale often show significant differences, implying that the scores given by people are incomparable due to differences in usage and calibration of each score (see Griffin & Brenner (2008); Poston (2008) and citations therein). Such differences in human scores are usually interpreted through simplifying modeling assumptions about how miscalibrations present in the data. These modeling assumptions include linear models with additive biases corresponding to rater identity (Bürkner & Vuorre, 2019; Paul, 2011; Barr et al., 2013), models with rater identity-based scale-and-shift biases (Paul, 2011; Roos et al., 2011), mixed-effects models, among others (Wang & Shah, 2019). However, research has shown that issues of human judgment calibration are often more complex, causing significant violations to these simplified assumptions (see Griffin & Brenner (2008) and citations therein). In this work, while making minimal assumptions on the nature of miscalibration in human judgments, we provide non-parametric metrics to quantify the consistency of raters in reflecting varying levels of severity. Traditional non-parametric metrics like Kendall’s $\tau$ and area under the PR or ROC curves do not capture well the responsiveness to severity. Using a real-world dataset, we surface insights from our proposed metrics into the feedback patterns of different rater groups.
@@ 3 Setup
We consider a general setup with two different rater populations that reflect two contrasting safety feedback paradigms (Rottger et al., 2022): crowd raters who indicate safety preferences of a diverse population, and trained raters who follow detailed guidelines to prescribe what is safe or unsafe.
@@
    Crowd raters provide pluralistic safety feedback using an ordinal scale, with each rating representing the perception of a certain rater group. Each crowd rater reviews the given item and provides an integer score on a 0-$K$ Likert scale, where 0 is not harmful and $K$ is completely harmful.
@@
    Trained raters strictly follow a set of prescribed rules. For each item, they provide a binary score of 0 or 1 indicating their feedback of safe or unsafe respectively.
@@ 3.1 Data model
To formalize our assumptions regarding the scores given by the individual raters, i.e., individual trained raters and individual crowd raters, we define a data model. Let $V_{i}$ be the true underlying severity of item $i$. Let $V^{j}_{i}$ be the true severity of item $i$ as perceived by rater $j$. Then,
$V^{j}_{i}=f_{j}(V_{i})$ (1)
where $f_{j}$ is the rater-specific perception function. This model acknowledges that true underlying severity $V$ of items is not directly measurable but is still the principal factor that monotonically influences every rater’s judgment. In the case of trained raters, $f_{j}$ represents the shared understanding of severity prescribed by their strict guidelines that serve as a framework to operationalize the theoretical notion of severity. In the case of crowd raters, $f_{j}$ represents some perception of true severity based on their lived experiences. In other words, perception of rater $j$ is primarily distributed along the dimension $V^{j}$ such that $V^{j}_{i}$ represents the unobservable true severity of item $i$ as perceived by rater $j$. Hereon, we use the notation $V^{j}$ for true severity as perceived by rater $j$ and the notation $V$ for the true underlying severity, aka simply, true severity.
The relationship between $V^{j}_{i}$ and the scores $S_{ij}$ of the two rater populations can then be stated via response functions as:
$M_{jk}(V^{j}_{i})\coloneq P(S_{ij}\geq k|V^{j}_{i})$ (2)
where $M_{jk}(V^{j}_{i})$ is non-decreasing in $V^{j}_{i}$. Further, we have that,
  • for trained raters, $S_{ij}\in\{0,1\}$, where $S_{ij}=1$ with probability $M_{j1}(V^{j}_{i})$.
  • for crowd raters, $S_{ij}\in\{0,1,\dots,K\}$, where $S_{ij}=k$ with probability $M_{jk}(V^{j}_{i})-M_{j(k+1)}(V^{j}_{i})$ and $M_{j(K+1)}(V^{j}_{i})=0$.
We assume $V$ to be unidimensional following the standard assumption of a unidimensional latent trait found in methods like Item Response Theory (Samejima, 1968). We note that the assumption limits our ability to simultaneously measure responsiveness with respect to multiple dimensions of severity.
@@ 3.2 Responsiveness to severity
Having a Likert scale allows crowd raters to express their safety feedback on a severity spectrum rather than classifying items as safe or unsafe. However, Likert scales themselves do not guarantee that rater scores will meaningfully reflect the severity of violations. For example, there may be raters who only use the ends of the scale, or those who cluster all their feedback around certain scores. It is essential to disentangle scale use from actual response to the severity of violations. While one might ideally want to define responsiveness as a direct relationship between scores $S$ of a rater $j$ that capture $V^{j}$ and true severity $V$, this is not practically feasible. True severity $V$ is a latent construct that cannot be directly or exactly determined. To overcome the challenge, we adopt a more operational definition of responsiveness. We formally define responsiveness to the severity of violations as being composed of the following two properties that can be quantified using observable data:
@@
  • Ability to stochastically order severity. If a rater is responsive to the severity of violations, then a higher score from them should correspond to a higher probability of true severity $V$ crossing any threshold $T=t$. This is the principle of first-order stochastic dominance, which can be stated as $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ for all $T\in\mathcal{T}$ when $s_{1}>s_{2}$.
@@
  • Ability to discriminate between distinct levels of severity. If a rater is responsive to the severity of violations, then they should be able to discriminate the items whose true severities $V$ are above a given threshold $T=t$ from those below that threshold. This means that $P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$ for all $T\in\mathcal{T}$.
@@
Intuitively, these two properties together signify that a rater who is responsive to the severity of violations is able to convey both granular and broader variations in the true underlying severity $V$ using the Likert scale. Our approach to characterizing responsiveness to severity relies on a notion of stochastic ordering that is related to the classic decision-theoretic concepts of stochastic ordering and outcome monotonicity (Birnbaum & Navarrete, 1998). The stochastic ordering property can further be related to monotonicity of a calibration curve or reliability diagram (DeGroot & Fienberg, 1983) in the case that we consider calibration of crowd rater scores against some binarized reference for severity. The property concerning the ability of raters to discriminate between distinct levels of severity is related to the notion of discriminability from signal detection theory (McNicol, 2005) used to motivate the design of metrics such as the area under the receiver operating characteristic (ROC) curve and Kendall’s $\tau$ (Kendall, 1938). Furthermore, the property is also related to the notion of discrimination between different levels of the latent trait in Item Response Theory (Samejima, 1968) and Mokken Scale Analysis (Mokken, 1971).
@@ 4 Metric design
Next, our aim is to quantify the responsiveness of the diverse crowd raters to the severity of violations in order to evaluate and compare them. Such a quantification can be achieved by individually quantifying the two properties that constitute responsiveness. We note that during safety alignment, the notion of true severity $V$ is operationalized depending on the choice of the feedback paradigm. When safety alignment is conducted based on feedback of trained raters, true severity $V$ is operationalized as captured by their guidelines, i.e., $V^{g}$. When safety alignment is conducted based on feedback of crowd raters, true severity $V$ is operationalized as captured by the collective judgment of the crowd, i.e., $V^{c}$. In the latter case, there is no dependence on the existence of any guidelines.
@@
Either way, true severity $V$ remains directly unobservable. We take $U$ to denote the observable binary reference such that $U=1$ signals $V>T$, i.e., unsafe. We can now reformulate the inequalities presented above for the ability to stochastically order and the ability to discriminate by leveraging $U=1$ in place of $V>T$ and $U=0$ in place of $V\leq T$. Note that since we do not assume a fixed threshold $T$, it is not straightforward to directly substitute $U=1$ into the threshold-specific inequalities. However, we show that if the two inequalities hold for all individual thresholds $T$, then they also hold for $U=1$. Proofs for both the inequalities are in appendix A.
  • Ability to stochastically order. We had that $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ when $s_{1}>s_{2}$, for all $T=t$. Hence, we can prove that $s_{1}>s_{2}$, $P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$.
  • Ability to discriminate. We had that $P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$ for all $T=t$. Hence, we can prove that $P(S\geq s|U=1)\geq P(S\geq s|U=0)$.
Depending on the operationalization of true severity $V$ under consideration, we can obtain the observable binary reference $U$, i.e., guideline-based reference or crowd-based reference, as described:
  1. 1.
    We can treat the binary scores of the trained raters as the $U$. We assign the individual binary scores of every trained rater to the items by replicating each item for every trained rater. When $U$ is obtained from trained raters, we are quantifying the responsiveness to severity $V^{g}$.
  2. 2.
    We can derive the $U$ from the crowd raters by excluding the rater(s) being evaluated to maintain the independence of $S$ and $T$ by construction. We binarize the individual Likert scores of the crowd raters, excluding those being evaluated, using scores $k\in[1,K]$ as boundaries. We assign these individual binary scores to the items by replicating each item for every crowd rater. When $U$ is obtained from crowd raters, we are quantifying the responsiveness to severity $V^{c}$.
The design of our metrics itself is agnostic to the choice of the reference, and we work with both.